In natural numbers, 4x + 5 = – 7 has                              solution. 
 In natural numbers, 4x + 5 = – 7 has                              solution. 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given integers, 4x + 5 =–7.
Transpose 5 to RHS then it becomes -5.

4x = - 7 - 5
4x = - 12
x = -12/4
x= - 3

In natural numbers, 4x + 5 =–7 has no solution.

Thanks